OpenStudy (angelwings996):

Find the 17th term of the sequence a_18 = 32, d=-4

OpenStudy (anonymous):

18a = 32 32 / 18 = a a = ?

OpenStudy (angelwings996):

1.78

OpenStudy (anonymous):

That solves for a. I've no clue what you mean by 17th term, as it's not a sequence, and I don't see the point to value d.

OpenStudy (angelwings996):

I'm not sure either, that's what the problem in the textbook is asking

OpenStudy (angelwings996):

It says in the back of the book that it is 36

OpenStudy (anonymous):

Quite a poorly worded question.

OpenStudy (angelwings996):

I had a problem earlier asking for the 32nd term for the sequence, but it actually gave me a list of numbers as this one doesn't

OpenStudy (anonymous):

general form for arithmetic sequence: \[a_{n}=a_{1}+(n-1)d\] since we know 18th term we can substitute: \[(32)=a_{1}+((18)-1)(-4)\] solve for a1

OpenStudy (angelwings996):

Uhmm...I got -2.125 ?

OpenStudy (angelwings996):

Did I do it wrong?

OpenStudy (anonymous):

\[32 = a_{1} - 68\] \[a_{1} = 100\]

OpenStudy (angelwings996):

Ohh, I wasn't paying attention and divided them instead of adding

OpenStudy (angelwings996):

@brusmack what do I do now?

OpenStudy (anonymous):

now you can fill in the general equation \[a_{n}=100+(n-1)(-4)\]

OpenStudy (anonymous):

use n=17 for the 17th term

OpenStudy (angelwings996):

Okay, so would it then be like \[a_{17}=100+(18-1)(-4)\]

OpenStudy (angelwings996):

Wait 17 not 18

OpenStudy (anonymous):

yep, so you should get a(17)=36

OpenStudy (angelwings996):

Okay, thank you so much!!

OpenStudy (anonymous):

no problem